# On local Galois deformation rings: generalised reductive groups

### v1

@@ 2.3. Extensions
Mapping $(N\times\Delta,\ast)$ to $(\omega^{\ast},c^{\ast})$ induces a bijection from the set of rigidified extensions of $N$ by $\Delta$ and the set of generalised $2$-cocycles of $\Delta$ with values in $N$.
@@
If $N$ is abelian then $\operatorname{Int}_{N}=\operatorname{id}_{N}$ and the result is well known. In general, the assertion follows from [49, Satz I], a nice exposition is given in [60, Section 2.2] and Lemma follows from [60, Lemma 2.11]. One could also deduce the assertion from [24]: property (3) implies that the element denoted by $f_{3}(x,y,z)$ in [24, Section 7] is $1_{N}$. Theorem 8.1 in [24] then implies that every generalised $2$-cocycle comes from a rigidified extension. ∎
@@
Since $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(\kappa),\mu_{p^{n}})=\kappa^{\times}/(\kappa^{\times})^{p^{n}}$ by [54, Tag 040Q] and $\kappa$ is a perfect field of characteristic $p$, we have $\mu(\kappa)=1$ and $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(\kappa),\mu)=0$. So we can see $c$ as a map to $H_{0}(\kappa)$, which we will denote by $\tilde{c}$. If $A$ is a $\kappa$-algebra we let $\tilde{c}_{A}$ be the composition of $c$ with the map $H_{0}(\kappa)\to H_{0}(A)$.
@@
Let $\tilde{\omega}:=L\circ\omega$, i.e. for all $\delta\in\Delta$, we have $\omega(\delta)\circ\pi=\pi\circ\tilde{\omega}(\delta)$. We claim that the pair $(\tilde{\omega},\tilde{c})$ satisfies the identities (C1), (C2), (C3) and (C4) above. Clearly the map $\tilde{c}$ satisfies (C4). Via the isomorphism $H^{0}(\kappa)\cong G^{0}(\kappa)$ the maps $\omega_{\kappa}$ and $\tilde{\omega}_{\kappa}$ can be identified, so (C3) holds for $(\tilde{\omega},\tilde{c})$. Property (C2) holds, as $L$ is a homomorphism. By applying $L$ to the identity (C1) for $G^{0}$, we get
@@ 3. Deformation rings
Let $\rho:\Gamma\rightarrow G(\kappa)$ be a continuous representation, where $G$ is a smooth affine $\mathcal{O}$-group scheme and $\kappa$ is either a finite extension of $k$, a finite extension of $L$ or a local field of characteristic $p$ containing $k$ equipped with natural topology. We will refer to such fields $\kappa$ as a finite or a local $\mathcal{O}$-field. In this section we will study the deformation theory of $\rho$. We first define a ring of coefficients $\Lambda$ over which the deformation problem is defined following [8, Section 3.5].
@@
    $A$ is a finite $\mathcal{O}/\varpi^{n}$-module for some $n\geq 1$ with the discrete topology on $A$;
@@
    $A$ is a finite dimensional $L$-vector space with the $p$-adic topology on $A$;
@@
    $A$ is a $\Lambda^{0}[1/t]/\varpi^{n}\Lambda^{0}[1/t]$-module of finite length for some $n\geq 1$ and we put the induced topology on $A$.
@@
which implies (13). The second part follows from local Euler–Poincaré characteristic formula and local Tate duality, see for example [9, Theorem 3.4.1]. ∎
@@
$1\rightarrow 1+\varepsilon\operatorname{Lie}G_{\kappa}\rightarrow G(\kappa[\varepsilon])\rightarrow G(\kappa)\rightarrow 1$
@@
The map $Z^{1}(\Gamma,\operatorname{ad}\rho)\rightarrow D^{\square}_{\rho,G}(\kappa[\varepsilon])$, $\Phi\mapsto[\gamma\mapsto\rho(\gamma)(1+\varepsilon\Phi(\gamma))]$ is an isomorphism of $\kappa$-vector spaces. Since $\Gamma$ satisfies Mazur’s $p$-finiteness condition [9, Theorem 3.4.1] implies that $h^{1}(\Gamma,\operatorname{ad}\rho)$ is finite. Lemma 3.1 implies that $Z^{1}(\Gamma,\operatorname{ad}\rho)$ is finite dimensional, thus $R^{\square}_{\rho,G}$ is noetherian. ∎
@@
If $G$ and $H$ are profinite, this is [50, Section I.1, Proposition 1]. In general, let $N$ be an open profinite subgroup of $G$. Then $N^{\prime}:=f(N)$ is an open profinite subgroup of $H$ and we find a continuous map $s:N^{\prime}\to N$ with $f|_{N}\circ s=\operatorname{id}_{N^{\prime}}$. Let $R\subseteq H$ be a set of $N^{\prime}$-coset representatives so that $H=\bigcup_{r\in R}rN^{\prime}$ and choose $g_{r}\in G$ with $f(g_{r})=r$ for each $r\in R$. We can extend $s$ to a continuous map $\tilde{s}:H\to G$ by defining $\tilde{s}(rn^{\prime}):=g_{r}s(n^{\prime})$ for all $n^{\prime}\in N^{\prime}$ and see, that $f\circ\tilde{s}=\operatorname{id}_{H}$, as desired. ∎
@@
Let $\kappa$ be either a finite or a local $\mathcal{O}$-field, let $\Lambda$ be a coefficient ring for $\kappa$ and let $(A,\mathfrak{m}_{A})\in\mathfrak{A}_{\Lambda}$. Then there is a module-finite $\Lambda^{0}$-subalgebra $R$ of $A$ with $R\otimes_{\Lambda^{0}}\Lambda=A$. In particular, $R$ is a profinite open subring of $A$.
@@
Part (2). Surjectivity of $\varphi$ follows from smoothness of $G$. Then $R^{\prime}:=f(R)$ is a compact subalgebra of $B$ with $R^{\prime}\otimes_{\Lambda^{0}}\Lambda=B$. Since the topology of $B$ is by definition induced by an arbitrary open $\Lambda^{0}$-lattice, we see that $R^{\prime}$ is open in $B$. Since $G$ is smooth, we have a surjection $G(R)\to G(R^{\prime})$ of open profinite subgroups of $G(A)$ and $G(B)$ respectively. Since any surjective continuous homomorphism between profinite groups is open [^1] Indeed, any continuous surjective homomorphism from a compact group onto a Hausdorff group is open by Remark 1 after Proposition 24 in [12, Ch. III, §2, no. 8]., we see that $\varphi$ is open. The assertion follows from 3.3. ∎
@@
If $\kappa$ is a finite field then the argument of [8, Proposition 4.3], where $G=\operatorname{GL}_{d}$, $H=\operatorname{GL}_{1}$, $\varphi=\det$, goes through verbatim. If $\kappa$ is a local field then one needs to additionally verify that obstruction to lifting can be realised by a continuous $2$-cocycle, as noted in [8, Remark 4.4]. The solution to this problem is explained in detail in [18, Lecture 6], when $\kappa$ is a local field of characteristic zero and $H$ is trivial. We sketch how these arguments fit together.
@@
$\widetilde{\phi}:\tilde{R}:=R_{\varphi\circ\rho}^{\square}\llbracket x_{1},\ldots,x_{r}\rrbracket\twoheadrightarrow R^{\square}_{\rho}.$
@@
The surjection of locally profinite groups $G(R_{u})\twoheadrightarrow G(R^{\square}_{\rho,G,i})$ has a continuous section by 3.5. Once we have this the proof is identical to the proof of [8, Proposition 4.3]. Moreover, if $\Gamma=\Gamma_{F}$ then (16) follows from Lemma 3.1. ∎
@@
Since $A$ is complete intersection we may write $A=S/(g_{1},\ldots,g_{s})$, where $S$ is a complete local regular ring and $g_{1},\ldots,g_{s}$ is a regular sequence in $S$. We choose $\tilde{f}_{1},\ldots,\tilde{f}_{t}\in S\llbracket x_{1},\ldots,x_{r}\rrbracket$ that map to $f_{1},\ldots,f_{t}$. Since when we quotient out by a relation the dimension either goes down by one or stays then same, the assumption on the dimension of $B$ implies that $\dim B=\dim A+r-t$ and $g_{1},\ldots,g_{s},\tilde{f}_{1},\ldots,\tilde{f}_{t}$ can be extended to a system of parameters of $S\llbracket x_{1},\ldots,x_{r}\rrbracket$, and hence is a regular sequence in the regular ring $S\llbracket x_{1},\ldots,x_{r}\rrbracket$. Thus $B$ is complete intersection and also $C:=S\llbracket x_{1},\ldots,x_{r}\rrbracket/(\tilde{f}_{1},\ldots,\tilde{f}_{t})$ is complete intersection. Since $k\otimes_{S}B=k\otimes_{A}B$, it follows from miracle flatness [41, Theorem 23.1] that $C$ is flat over $S$, and hence $B=C/(g_{1},\ldots,g_{s})$ is flat over $A=S/(g_{1},\ldots,g_{s})$. ∎
@@
The proof is the same as the proof of [9, Theorem 3.3.1] for $\operatorname{GL}_{d}$, which in turn is based on [35, Proposition 9.5], where the case $G=\operatorname{GL}_{d}$ and $\mathrm{char}(\kappa)=0$ is treated. ∎
@@ 3.1. Central extensions
has kernel $Z\cap G^{\prime}$, so it is injective. If $([g],h)\in G(A)/G^{\prime}(A)\times_{H(A)/H^{\prime}(A)}H(A)$ then, since $G\rightarrow H$ is surjective, there exists a finite flat extension $A\hookrightarrow B$ and $\tilde{g}\in G(B)$, which maps to $h\in H(A)\subseteq H(B)$. Then $\tilde{g}g^{-1}\in H^{\prime}(B)=G^{\prime}(B)$, so $[\tilde{g}]=[g]$ in $G(B)/G^{\prime}(B)$. By sheafification this implies, that (24) is bijective.
@@
Let $A^{\prime}=(R^{\square}_{H/H^{\prime}}/\varpi)^{\mathrm{red}}$, $B^{\prime}=(R^{\square}_{G/G^{\prime}}/\varpi)^{\mathrm{red}}$, $A=(R^{\square}_{H}/\varpi)^{\mathrm{red}}$, where the superscript $\mathrm{red}$ indicates reduced rings. We have shown in [43, Corollary 9.9] that $A^{\prime}\rightarrow B^{\prime}$ is flat and the dimension $d$ of the fibre $\kappa\otimes_{A^{\prime}}B^{\prime}$ is equal to
@@
Hence, $A\rightarrow B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A$ is flat and the fibre has dimension $d$. Thus $\dim A+d=\dim B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A$ by [54, Tag 00ON]. 3.13 implies that $(B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A)^{\mathrm{red}}\cong(R^{\square}_{G}/\varpi)^{\mathrm{red}}$. Hence, $\dim R^{\square}_{G}/\varpi=\dim A+d=\dim R^{\square}_{H}/\varpi+d$. ∎
@@
Let $e\geq 0$ be an integer such that $p^{e}\cdot\mathrm{m}=0$. Let $A=R^{\square}_{H}/\varpi$ and let $\mathfrak{m}$ be the maximal ideal of $A$. We choose elements $a_{1},\ldots,a_{r}\in\mathfrak{m}$ such that their images build a basis of $\mathfrak{m}/\mathfrak{m}^{2}$ as a $\kappa$-vector space. Let $B=A[x_{1},\ldots,x_{n}]/I$, where $I$ is the ideal generated by $x_{i}^{p^{e}}-a_{i}$ for $1\leq i\leq r$. Then $B$ is is finite flat over $A$ and every $x\in 1+\mathfrak{m}$ is a $p^{e}$-th power in $B^{\times}$. Let $A^{\mathrm{red}}$ and $B^{\mathrm{red}}$ be the maximal reduced quotients of $A$ and $B$, respectively. Then $A^{\mathrm{red}}$ is a subring of $B^{\mathrm{red}}$ and every $x\in 1+\mathfrak{m}_{A^{\mathrm{red}}}$ is a $p^{e}$-th power in $B^{\mathrm{red}}$. Let $\rho^{\prime}:\Gamma\rightarrow H(A^{\mathrm{red}})$ be the specialisation of the universal deformation $\Gamma\rightarrow H(R^{\square}_{H})$ along $R^{\square}_{H}\rightarrow A^{\mathrm{red}}$.
Since $Z$ is a finite product of $\mu_{p^{n}}$ with $n\leq e$ and $B^{\mathrm{red}}$ is reduced and of characteristic $p$, we have $Z(B^{\mathrm{red}})=1$. Since $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(A),\mu_{p^{n}})=A^{\times}/(A^{\times})^{p^{n}}$ by [54, Tag 040Q], it follows from (25) that there exists a continuous representation $\tilde{\rho}:\Gamma\rightarrow G(B^{\mathrm{red}})$ deforming $\rho$ such that $\varphi\circ\tilde{\rho}=\rho^{\prime}\otimes_{A^{\mathrm{red}}}B^{\mathrm{red}}$. This induces a map of $A$-algebras $R^{\square}_{G}/\varpi\rightarrow B^{\mathrm{red}}$. This map is finite as $B$ is a finite $A$-algebra. Hence $\dim(R^{\square}_{G}/\varpi)\geq\dim B^{\mathrm{red}}=\dim B=\dim A$, where the last equality follows as $B$ is finite and flat over $A$. Lemma 3.15 implies that $\dim(R^{\square}_{G}/\varpi)\leq\dim A$. ∎

### v2

@@ 2.3. Extensions
Mapping $(N\times\Delta,\ast)$ to $(\omega^{\ast},c^{\ast})$ induces a bijection from the set of rigidified extensions of $N$ by $\Delta$ to the set of generalised $2$-cocycles of $\Delta$ with values in $N$.
@@
If $N$ is abelian then $\operatorname{Int}_{N}=\operatorname{id}_{N}$ and the result is well known. In general, the assertion follows from [54, Satz I], a nice exposition is given in [65, Section 2.2] and Lemma follows from [65, Lemma 2.11]. One could also deduce the assertion from [24]: property (3) implies that the element denoted by $f_{3}(x,y,z)$ in [24, Section 7] is $1_{N}$. Theorem 8.1 in [24] then implies that every generalised $2$-cocycle comes from a rigidified extension. ∎
@@
Since $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(\kappa),\mu_{p^{n}})=\kappa^{\times}/(\kappa^{\times})^{p^{n}}$ by [59, Tag 040Q] and $\kappa$ is a perfect field of characteristic $p$, we have $\mu(\kappa)=1$ and $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(\kappa),\mu)=0$. So we can see $c$ as a map to $H_{0}(\kappa)$, which we will denote by $\tilde{c}$. If $A$ is a $\kappa$-algebra we let $\tilde{c}_{A}$ be the composition of $c$ with the map $H_{0}(\kappa)\to H_{0}(A)$.
@@
[…]
@@ 3. Deformation rings
Let $\rho:\Gamma\rightarrow G(\kappa)$ be a continuous representation, where $G$ is a smooth affine $\mathcal{O}$-group scheme and $\kappa$ is either a finite extension of $k$, a finite extension of $L$ or a local field of characteristic $p$ containing $k$ equipped with natural topology. We will refer to such fields $\kappa$ as a finite or a local $\mathcal{O}$-field. In this section we will study the deformation theory of $\rho$. We first define a ring of coefficients $\Lambda$ over which the deformation problem is defined following [6, Section 3.5].
@@
    AA is a finite $\mathcal{O}/\varpi^{n}$-module for some $n\geq 1$ with the discrete topology on $A$;
@@
    AA is a finite dimensional $L$-vector space with the $p$-adic topology on $A$;
@@
    AA is a $\Lambda^{0}[1/t]/\varpi^{n}\Lambda^{0}[1/t]$-module of finite length for some $n\geq 1$ and we put the induced topology on $A$.
@@
which implies (13). The second part follows from local Euler–Poincaré characteristic formula and local Tate duality, see for example [7, Theorem 3.4.1]. ∎
@@
$0\rightarrow\operatorname{Lie}G_{\kappa}\rightarrow G(\kappa[\varepsilon])\rightarrow G(\kappa)\rightarrow 0.$
@@
The map $Z^{1}(\Gamma,\operatorname{ad}\rho)\rightarrow D^{\square}_{\rho,G}(\kappa[\varepsilon])$, $\Phi\mapsto[\gamma\mapsto\rho(\gamma)(1+\varepsilon\Phi(\gamma))]$ is an isomorphism of $\kappa$-vector spaces. Since $\Gamma$ satisfies Mazur’s $p$-finiteness condition [7, Theorem 3.4.1] implies that $h^{1}(\Gamma,\operatorname{ad}\rho)$ is finite. Lemma 3.1 implies that $Z^{1}(\Gamma,\operatorname{ad}\rho)$ is finite dimensional, thus $R^{\square}_{\rho,G}$ is noetherian. ∎
@@
If $G$ and $H$ are profinite, this is [55, Section I.1, Proposition 1]. In general, let $N$ be an open profinite subgroup of $G$. Then $N^{\prime}:=f(N)$ is an open profinite subgroup of $H$ and we find a continuous map $s:N^{\prime}\to N$ with $f|_{N}\circ s=\operatorname{id}_{N^{\prime}}$. Let $R\subseteq H$ be a set of $N^{\prime}$-coset representatives so that $H=\bigcup_{r\in R}rN^{\prime}$ and choose $g_{r}\in G$ with $f(g_{r})=r$ for each $r\in R$. We can extend $s$ to a continuous map $\tilde{s}:H\to G$ by defining $\tilde{s}(rn^{\prime}):=g_{r}s(n^{\prime})$ for all $n^{\prime}\in N^{\prime}$ and see, that $f\circ\tilde{s}=\operatorname{id}_{H}$, as desired. ∎
@@
Let $\kappa$ be a local $\mathcal{O}$-field, let $\Lambda$ be a coefficient ring for $\kappa$ and let $(A,\mathfrak{m}_{A})\in\mathfrak{A}_{\Lambda}$. Then there is a module-finite $\Lambda^{0}$-subalgebra $R$ of $A$ with $R\otimes_{\Lambda^{0}}\Lambda=A$. In particular, $R$ is a profinite open subring of $A$.
@@
Part (2). Surjectivity of $\varphi$ follows from smoothness of $G$. Then $R^{\prime}:=f(R)$ is a compact subalgebra of $B$ with $R^{\prime}\otimes_{\Lambda^{0}}\Lambda=B$. Since the topology of $B$ is by definition induced by an arbitrary open $\Lambda^{0}$-lattice, we see that $R^{\prime}$ is open in $B$. Since $G$ is smooth, we have a surjection $G(R)\to G(R^{\prime})$ of open profinite subgroups of $G(A)$ and $G(B)$ respectively. Since any surjective continuous homomorphism between profinite groups is open [^2] Indeed, any continuous surjective homomorphism from a compact group onto a Hausdorff group is open by Remark 1 after Proposition 24 in [11, Ch. III, §2, no. 8]., we see that $\varphi$ is open. The assertion follows from 3.3. ∎
@@
If $\kappa$ is a finite field then the argument of [6, Proposition 4.3], where $G=\operatorname{GL}_{d}$, $H=\operatorname{GL}_{1}$, $\varphi=\det$, goes through verbatim. If $\kappa$ is a local field then one needs to additionally verify that obstruction to lifting can be realised by a continuous $2$-cocycle, as noted in [6, Remark 4.4]. The solution to this problem is explained in detail in [18, Lecture 6], when $\kappa$ is a local field of characteristic zero and $H$ is trivial. We sketch how these arguments fit together.
@@
$\tilde{\phi}:\tilde{R}:=R_{\varphi\circ\rho}^{\square}\llbracket x_{1},\ldots,x_{r}\rrbracket\twoheadrightarrow R^{\square}_{\rho}.$
@@
The surjection of locally profinite groups $G(R_{u})\twoheadrightarrow G(R^{\square}_{\rho,G,i})$ has a continuous section by 3.5. Once we have this the proof is identical to the proof of [6, Proposition 4.3]. Moreover, if $\Gamma=\Gamma_{F}$ then (16) follows from Lemma 3.1. ∎
@@
Since $A$ is complete intersection we may write $A=S/(g_{1},\ldots,g_{s})$, where $S$ is a complete local regular ring and $g_{1},\ldots,g_{s}$ is a regular sequence in $S$. We choose $\tilde{f}_{1},\ldots,\tilde{f}_{t}\in S\llbracket x_{1},\ldots,x_{r}\rrbracket$ that map to $f_{1},\ldots,f_{t}$. Since when we quotient out by a relation the dimension either goes down by one or stays then same, the assumption on the dimension of $B$ implies that $\dim B=\dim A+r-t$ and $g_{1},\ldots,g_{s},\tilde{f}_{1},\ldots,\tilde{f}_{t}$ can be extended to a system of parameters of $S\llbracket x_{1},\ldots,x_{r}\rrbracket$, and hence is a regular sequence in the regular ring $S\llbracket x_{1},\ldots,x_{r}\rrbracket$. Thus $B$ is complete intersection and also $C:=S\llbracket x_{1},\ldots,x_{r}\rrbracket/(\tilde{f}_{1},\ldots,\tilde{f}_{t})$ is complete intersection. Since $k\otimes_{S}B=k\otimes_{A}B$, it follows from miracle flatness [43, Theorem 23.1] that $C$ is flat over $S$, and hence $B=C/(g_{1},\ldots,g_{s})$ is flat over $A=S/(g_{1},\ldots,g_{s})$. ∎
@@
The proof is the same as the proof of [7, Theorem 3.3.1] for $\operatorname{GL}_{d}$, which in turn is based on [36, Proposition 9.5], where the case $G=\operatorname{GL}_{d}$ and $\mathrm{char}(\kappa)=0$ is treated. ∎
@@ 3.1. Central extensions
has kernel $Z\cap G^{\prime}$, so it is injective. If $([g],h)\in G(A)/G^{\prime}(A)\times_{H(A)/H^{\prime}(A)}H(A)$ then, since $G\rightarrow H$ is surjective, there exists a finite flat extension $A\hookrightarrow B$ and $\tilde{g}\in G(B)$, which maps to $h\in H(A)\subseteq H(B)$. Then $\tilde{g}g^{-1}\in H^{\prime}(B)=G^{\prime}(B)$, so $\tilde{g}]=[g]$ in $G(B)/G^{\prime}(B)$. By sheafification this implies, that (24) is bijective.
@@
Let $A^{\prime}=(R^{\square}_{H/H^{\prime}}/\varpi)^{\mathrm{red}}$, $B^{\prime}=(R^{\square}_{G/G^{\prime}}/\varpi)^{\mathrm{red}}$, $A=(R^{\square}_{H}/\varpi)^{\mathrm{red}}$, where the superscript $\mathrm{red}$ indicates reduced rings. We have shown in [48, Corollary 9.9] that $A^{\prime}\rightarrow B^{\prime}$ is flat and the dimension $d$ of the fibre $\kappa\otimes_{A^{\prime}}B^{\prime}$ is equal to
@@
Hence, $A\rightarrow B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A$ is flat and the fibre has dimension $d$. Thus $\dim A+d=\dim B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A$ by [59, Tag 00ON]. 3.13 implies that $(B^{\prime}\operatorname{\widehat{\otimes}}_{A^{\prime}}A)^{\mathrm{red}}\cong(R^{\square}_{G}/\varpi)^{\mathrm{red}}$. Hence, $\dim R^{\square}_{G}/\varpi=\dim A+d=\dim R^{\square}_{H}/\varpi+d$. ∎
@@
Let $e\geq 0$ be an integer such that $p^{e}\cdot\mathrm{m}=0$. Let $A=R^{\square}_{H}/\varpi$ and let $\mathfrak{m}$ be the maximal ideal of $A$. We choose elements $a_{1},\ldots,a_{r}\in\mathfrak{m}$ such that their images form a basis of $\mathfrak{m}/\mathfrak{m}^{2}$ as a $\kappa$-vector space. Let $B=A[x_{1},\ldots,x_{n}]/I$, where $I$ is the ideal generated by $x_{i}^{p^{e}}-a_{i}$ for $1\leq i\leq r$. Then $B$ is finite flat over $A$ and every $x\in 1+\mathfrak{m}$ is a $p^{e}$-th power in $B^{\times}$. Let $A^{\mathrm{red}}$ and $B^{\mathrm{red}}$ be the maximal reduced quotients of $A$ and $B$, respectively. Then $A^{\mathrm{red}}$ is a subring of $B^{\mathrm{red}}$ and every $x\in 1+\mathfrak{m}_{A^{\mathrm{red}}}$ is a $p^{e}$-th power in $B^{\mathrm{red}}$. Let $\rho^{\prime}:\Gamma\rightarrow H(A^{\mathrm{red}})$ be the specialisation of the universal deformation $\Gamma\rightarrow H(R^{\square}_{H})$ along $R^{\square}_{H}\rightarrow A^{\mathrm{red}}$.
Since $Z$ is a finite product of $\mu_{p^{n}}$ with $n\leq e$ and $B^{\mathrm{red}}$ is reduced and of characteristic $p$, we have $Z(B^{\mathrm{red}})=1$. Since $H^{1}_{\mathrm{fppf}}(\operatorname{Spec}(A),\mu_{p^{n}})=A^{\times}/(A^{\times})^{p^{n}}$ by [59, Tag 040Q], it follows from (25) that there exists a continuous representation $\tilde{\rho}:\Gamma\rightarrow G(B^{\mathrm{red}})$ deforming $\rho$ such that $\varphi\circ\tilde{\rho}=\rho^{\prime}\otimes_{A^{\mathrm{red}}}B^{\mathrm{red}}$. This induces a map of $A$-algebras $R^{\square}_{G}/\varpi\rightarrow B^{\mathrm{red}}$. This map is finite as $B$ is a finite $A$-algebra. Hence $\dim(R^{\square}_{G}/\varpi)\geq\dim B^{\mathrm{red}}=\dim B=\dim A$, where the last equality follows as $B$ is finite and flat over $A$. Lemma 3.15 implies that $\dim(R^{\square}_{G}/\varpi)\leq\dim A$. ∎